OpenStudy (anonymous):

Blake has a total of 64 DVD's and CD's. If the number of CD's is eight more than three times the number of DVD's, how many DVD's does he have?

OpenStudy (anonymous):

Let A be the number of DVD's, B be the number of CD's. From the problem, we have two equations: A + B = 64 3A + 8 = B Substitute B in the first equation, we are left with: A + 3A + 8 = 64 -> 4A = 64 - 8 = 56 A = 56/4 = 14. So, he has 14 DVD's